# Robustly overfitting latents for flexible neural image compression

### v1

@@ Abstract
Neural image compression has made a great deal of progress. State-of-the-art models are based on variational autoencoders and are outperforming classical models. Neural compression models learn to encode an image into a quantized latent representation that can be efficiently sent to the decoder, which decodes the quantized latent into a reconstructed image. While these models have proven successful in practice, they lead to sub-optimal results due to imperfect optimization and limitations in the encoder and decoder capacity. Recent work shows how to use stochastic Gumbel annealing (SGA) to refine the latents of pre-trained neural image compression models. We extend this idea by introducing SGA+, which contains three different methods that build upon SGA. Further, we give a detailed analysis of our proposed methods, show how they improve performance, and show that they are less sensitive to hyperparameter choices. Besides, we show how each method can be extended to three- instead of two-class rounding. Finally, we show how refinement of the latents with our best-performing method improves the compression performance on the Tecnick dataset and how it can be deployed to partly move along the rate-distortion curve.
@@ 1 Introduction
Image compression allows efficient sending of an image between systems by reducing their size. There are two types of compression: lossless and lossy. Lossless image compression sends images perfectly without losing any quality and can thus be restored in their original format, such as the PNG format. Lossy compression, such as BPG (Bellard, 2014), JPEG (Wallace, 1992) or JPEG2000 (Skodras et al., 2001), loses some quality of the compressed image. Lossy compression aims to preserve as much of the quality of the reconstructed image as possible, compared to its original format, while allowing a significantly larger reduction in required storage.
@@
Traditional methods (Wallace, 1992; Skodras et al., 2001), especially lossless methods, can lead to limited compression ratios or degradation in quality. With the rise of deep learning, neural image compression is becoming a popular method (Theis et al., 2017; Toderici et al., 2017). In contrast with traditional methods, neural image compression methods have been shown to achieve higher compression ratios and less degradation in image quality (Ballé et al., 2018; Minnen et al., 2018; Lee et al., 2019). Additionally, neural compression techniques have shown improvements compared to traditional codecs for other data domains, such as video. (Agustsson et al., 2020; Habibian et al., 2019; Lu et al., 2019).
@@
In practice, neural lossy compression methods have proven to be successful and achieve state-of-the-art performance (Ballé et al., 2018; Minnen et al., 2018; Lee et al., 2019). These models are frequently based on variational autoencoders (VAEs) with an encoder-decoder structure (Kingma & Welling, 2013). The models are trained to minimize the expected rate-distortion (R-D) cost: $R+\lambda D$. Intuitively, one learns a mapping that encodes an image into a compressible latent representation. The latent representation is sent to a decoder and is decoded into a reconstructed image. The aim is to train the compression model in such way that it finds a latent representation that represents the best trade-off between the length of the bitstream for an image and the quality of the reconstructed image. Even though these models have proven to be successful in practice, they do have limited capacity when it comes to optimization and generalization. For example, the encoder’s capacity is limited which makes the latent representation sub-optimal (Cremer et al., 2018). Recent work (Campos et al., 2019; Guo et al., 2020; Yang et al., 2020) proposes procedures to refine the encoder or latents, which lead to better compression performance. Furthermore, in neural video compression, other work focuses on adapting the encoder (Aytekin et al., 2018; Lu et al., 2020) or finetuning a full compression model after training to improve the video compression performance (van Rozendaal et al., 2021).
@@
The advantage of refining latents (Campos et al., 2019; Yang et al., 2020) is that improved compression results per image are achieved while the model does not need to be modified. Instead, the latent representations for each individual image undergo a refining procedure. This results in a latent representation that obtains an improved bitstream and image quality over its original state from the pre-trained model. As mentioned in (Yang et al., 2020), the refining procedure for stochastic rounding with Stochastic Gradient Gumbel Annealing (SGA) considerably improves performance. In this paper, we introduce SGA+, an extension of SGA that further improves compression performance and is less sensitive to hyperparameter choices. The main contributions are:
  • We show how changing the probability space with more natural methods instead of SGA boosts the compression performance.
  • We propose the sigmoid scaled logit (SSL), which can smoothly interpolate between the approximate $\operatorname{atanh}$, linear, cosine and round. And we show how this function finds even better compression performance.
  • We demonstrate a generalization to rounding to three classes, that contains the two classes as a special case.
@@
Further, we show how SSL outperforms baselines on the Kodak dataset in terms of true loss curves. We show how our method generalizes to the Tecnick dataset in terms of peak signal-to-noise ratio (PSNR) versus the bits per pixel (BPP) in an R-D plot. In addition, we analyze the stability of all functions and show the effect of interpolation between different methods with SSL. Lastly, we propose a refining procedure at compression time that allows moving along the R-D curve when refining the latents with another $\lambda$ than a pre-trained model is trained on. The code of our proposed methods will be published at a later time on GitHub.
@@ 2 Preliminaries and related work
In lossy compression, the aim is to find a mapping of image $x$ where the distortion of the reconstructed image $\hat{x}$ is as little as possible compared to the original one while using as little storage as possible. Therefore, training a lossy neural image compression model presents a trade-off between minimizing the length of the bitstream for an image and minimizing the distortion of the reconstructed image (Ballé et al., 2017; Lee et al., 2019; Minnen et al., 2018; Theis et al., 2017).
Neural image compression models from (Ballé et al., 2017; Minnen et al., 2018; Theis et al., 2017), also known as hyperpriors, accomplish this kind of mapping with latent variables. An image $x$ is encoded onto a latent representation $y=g_{a}(x)$, where $g_{a}(\cdot)$ is the encoder. Next, $y$ is quantized $Q(y)=\hat{y}$ into a discrete variable that is sent losslessly to the decoder. The reconstructed image is given by: $\hat{x}=g_{s}(\hat{y})$, where $g_{s}(\cdot)$ represents the decoder. The rate-distortion objective that needs to be minimized for this specific problem is given by:
@@ 2.1 Latent optimization
Neural image compression models have been trained over a huge set of images to find an optimal encoding. Yet, due to difficulties in optimization or due to constraints on the model capacity, model performance is sub-optimal. To overcome these issues, another type of optimizing compression performance is proposed in (Campos et al., 2019; Yang et al., 2020) where they show how to find better compression results by utilizing pre-trained networks and keeping the encoder and decoder fixed but only adapting the latents. In these methods, a latent variable $y$ is iteratively adapted using differentiable operations at test time. The aim is to find a more optimal discrete latent representation $\hat{y}$. Therefore, the following minimization problem needs to be solved for an image $x$:
@@ 2.2 Stochastic Gumbel Annealing
$\log(\pi)=(-\operatorname{atanh}(v_{L})/\tau,-\operatorname{atanh}(v_{R})/\tau).$ (3)
Next, samples are drawn: $\mathbf{y}\sim\text{Gumbel-Softmax}(\pi,\tau)$ (Jang et al., 2016) and are multiplied and summed with the vector $\mathbf{v}_{r}$ to obtain the quantized representation: $\hat{v}=\sum_{i}{(v_{r,i}*y_{i})}$. As SGA aids the discretization gap, this method may not have optimal performance and may not be as robust to changes in its temperature rate $\tau$.
Besides SGA, (Yang et al., 2020) propose deterministic annealing (Agustsson et al., 2017), which follows almost the same procedure as SGA, but instead of sampling stochastically from the Gumbel Softmax, this method uses a deterministic approach by computing probabilities with the Softmax from $\log(\pi)$. In practice, this method has been shown to suffer from unstable optimization behavior.
@@ 2.3 Other methods
While methods such as SGA aim to optimize the latent variables for neural image compression at inference time, other approaches have been explored in recent research. (Guo et al., 2021) proposed a soft-then-hard strategy alongside a learned scaling factor for the uniform noise to achieve better compression and a smoother latent. These methods are used to fine-tune network parameters but not the latents directly. (Zhu et al., 2022) proposed using Swin-transformer-based coding instead of ConvNet-based coding. They showed that these transforms can achieve better compression with fewer parameters and shorter decoding times. (van Rozendaal et al., 2021) proposed to also fine-tune the decoder alongside the latent for video compression. While accommodating the additional cost of saving the model update, they demonstrated a gain of $\sim 1dB$. (Zhang et al., 2021) and (Dupont et al., 2021) proposed using implicit neural representations for video and image compression, respectively. (He et al., 2022) proposed an improved context model (SCCTX) and a change to the main transform (ELIC) that achieve strong compression results together. (El-Nouby et al., 2023) revisited vector quantization for neural image compression and demonstrated it performs on par with hyperprior-based methods. While these approaches change the training process, our work differs in that we only consider the inference process.
@@ 3 Methods
As literature has shown, refining the latents of pre-trained compression models with SGA leads to improved compression performance (Yang et al., 2020). In this section, we extend SGA by introducing SGA+ containing three other methods for the computation of the unnormalized log probabilities $\log(\pi)$ to overcome issues from its predecessor. We show how these methods behave in probability space. Furthermore, we show how the methods can be extended to three-class rounding.
(a) The probability space for different functions of SGA+: linear, cosine and SSL ($a=\frac{4}{3}$), along with $\operatorname{atanh}$. Solid lines denote the probability of flooring $\lfloor v\rfloor$ and dotted lines the probability of ceiling $\lceil v\rceil$.
(b) Three-class rounding for the extended version of linear. Solid lines denote the two-class rounding with $r=1$ and $n=1$. Dashed lines denote three-class rounding with $r=0.9$ and $n=1$ and the dotted lines indicate the smoothness for $r=1$ and $n=3$.
@@ 3.1 Two-class rounding
Recall from SGA that a variable $v$ is quantized to indicate the rounding direction to two classes and is centered between (0,1). Computation of the unnormalized log probabilities is obtained with $\operatorname{atanh}$ from Equation (3). When looking at the probability space from this function, see Figure 0(a), the $\operatorname{atanh}$ function can lead to sub-optimal performance when used to determine rounding probabilities. The problem is that gradients tend to infinity when the function approaches the limits of 00 and $1$. This is not ideal, as these limits are usually achieved when the discretization gap is minimal. In addition, the gradients may become larger towards the end of optimization. Further analyzing the probability space, we find that there are a lot of possibilities in choosing probabilities for rounding to two classes. However, there are some constraints: the probabilities need to be monotonic functions, and the probabilities for rounding down (flooring) and up (ceiling) need to sum up to one. Therefore, we introduce SGA+ and propose three methods that satisfy the above constraints and can be used to overcome the sub-optimality that the $\operatorname{atanh}$ function suffers from.
@@
where $a$ is the factor determining the shape of the function. SSL resembles exactly the linear for $a=1$. For $a=1.6$ and $a=0.65$ SSL roughly resembles the cosine and $\operatorname{atanh}$. For $a\rightarrow\infty$ the function tends to shape to (reversed) rounding.
@@ 3.2 Three-class rounding
The probability that $v$ is rounded is denoted by: $p(y=\lfloor v\rceil)\propto f_{3c}(w|r,n)$, where $w=v-\lfloor v\rceil$ is centered around zero. Further, we define the probability that $v$ is rounded $+1$ and rounded $-1$ is respectively given by: $p(y=\lfloor v\rceil-1)\propto f_{3c}(w-1|r,n)$ and $p(y=\lfloor v\rceil+1)\propto f_{3c}(w+1|r,n)$. The general notation for the three-class functions is given by:
@@
where $\operatorname{clip}(\cdot)$ clips the value at 00 and $1$, $r$ is the factor determining the height and steepness of the function and power $n$ controls the peakedness of the function.
@@
In Figure 0(b), the three-class rounding for the extension of Equation (5) can be found. As can be seen, the solid lines denote the special case of two-class rounding with $r=1$ and $n=1$, the dashed lines denote three-class rounding with $r=0.9$ and $n=1$ and the dotted lines denote the two-class rounding with $r=1$ and $n=3$, which shows a less peaked function. We can now compare an example of two- versus three-class rounding. Consider the case where we have variable $v=-0.95$. For two-class rounding there is only the chance of rounding to $-1$ with $p(y=\lfloor v\rceil)$ (red solid line), a chance to round to 00 with $p(y=\lfloor v\rceil+1)$ (green solid line) and zero chance to round to $-2$ with $p(y=\lfloor v\rceil-1)$ (yellow solid line). Now for three-class rounding, with $r=0.9$ and $n=1$, when $v=-0.95$ we find that there is a high chance to round to $-1$ with $p(y=\lfloor v\rceil)$ (red dashed line) and a small chance to round to 00 with $p(y=\lfloor v\rceil+1)$ (green dashed line) and a tiny chance to round to $-2$ with $p(y=\lfloor v\rceil-1)$ (yellow dashed line).

### v3

@@ Abstract
Neural image compression has made a great deal of progress. State-of-the-art models are based on variational autoencoders and are outperforming classical models. Neural compression models learn to encode an image into a quantized latent representation that can be efficiently sent to the decoder, which decodes the quantized latent into a reconstructed image. While these models have proven successful in practice, they lead to sub-optimal results due to imperfect optimization and limitations in the encoder and decoder capacity. Recent work shows how to use stochastic Gumbel annealing (SGA) to refine the latents of pre-trained neural image compression models. We extend this idea by introducing SGA+, which contains three different methods that build upon SGA. We show how our method improves the overall compression performance in terms of the R-D trade-off, compared to its predecessors. Additionally, we show how refinement of the latents with our best-performing method improves the compression performance on both the Tecnick and CLIC dataset. Our method is deployed for a pre-trained hyperprior and for a more flexible model. Further, we give a detailed analysis of our proposed methods and show that they are less sensitive to hyperparameter choices. Finally, we show how each method can be extended to three- instead of two-class rounding.
@@ 1 Introduction
Image compression allows efficient sending of an image between systems by reducing their size. There are two types of compression: lossless and lossy. Lossless image compression sends images perfectly without losing any quality and can thus be restored in their original format, such as the PNG format. Lossy compression, such as BPG Bellard (2014), JPEG Wallace (1992) or JPEG2000 Skodras et al. (2001), loses some quality of the compressed image. Lossy compression aims to preserve as much of the quality of the reconstructed image as possible, compared to its original format, while allowing a significantly larger reduction in required storage.
@@
Traditional methods Wallace (1992); Skodras et al. (2001), especially lossless methods, can lead to limited compression ratios or degradation in quality. With the rise of deep learning, neural image compression is becoming a popular method Theis et al. (2017); Toderici et al. (2017). In contrast with traditional methods, neural image compression methods have been shown to achieve higher compression ratios and less degradation in image quality Ballé et al. (2018); Minnen et al. (2018); Lee et al. (2019). Additionally, neural compression techniques have shown improvements compared to traditional codecs for other data domains, such as video. Agustsson et al. (2020); Habibian et al. (2019); Lu et al. (2019).
@@
In practice, neural lossy compression methods have proven to be successful and achieve state-of-the-art performance Ballé et al. (2018); Cheng et al. (2020); Minnen et al. (2018); Lee et al. (2019). These models are frequently based on variational autoencoders (VAEs) with an encoder-decoder structure Kingma and Welling (2013). The models are trained to minimize the expected rate-distortion (R-D) cost: $R+\lambda D$. Intuitively, one learns a mapping that encodes an image into a compressible latent representation. The latent representation is sent to a decoder and is decoded into a reconstructed image. The aim is to train the compression model in such way that it finds a latent representation that represents the best trade-off between the length of the bitstream for an image and the quality of the reconstructed image. Even though these models have proven to be successful in practice, they do have limited capacity when it comes to optimization and generalization. For example, the encoder’s capacity is limited which makes the latent representation sub-optimal Cremer et al. (2018). Recent work Campos et al. (2019); Guo et al. (2020); Yang et al. (2020) proposes procedures to refine the encoder or latents, which lead to better compression performance. Furthermore, in neural video compression, other work focuses on adapting the encoder Aytekin et al. (2018); Lu et al. (2020) or finetuning a full compression model after training to improve the video compression performance van Rozendaal et al. (2021).
@@
The advantage of refining latents Campos et al. (2019); Yang et al. (2020) is that improved compression results per image are achieved while the model does not need to be modified. Instead, the latent representations for each individual image undergo a refining procedure. This results in a latent representation that obtains an improved bitstream and image quality over its original state from the pre-trained model. As mentioned in Yang et al. (2020), the refining procedure for stochastic rounding with Stochastic Gradient Gumbel Annealing (SGA) considerably improves performance.
@@
In this paper, we introduce SGA+, an extension of SGA that further improves compression performance and is less sensitive to hyperparameter choices. The main contributions are: (i) showing how changing the probability space with more natural methods instead of SGA boosts the compression performance, (ii) proposing the sigmoid scaled logit (SSL), which can smoothly interpolate between the approximate $\operatorname{atanh}$, linear, cosine and round, (iii) demonstrating a generalization to rounding to three classes, that contains the two classes as a special case, and (iv) showing that SGA+ not only outperforms SGA on a similar pre-trained mean-scale hyperprior model as in Yang et al. (2020), but also achieves an even better performance for the pre-trained models of Cheng et al. (2020).  Further, we show how SSL outperforms baselines in an R-D plot on the Kodak dataset, in terms of peak signal-to-noise ratio (PSNR) versus the bits per pixel (BPP) and in terms of true loss curves. Additionally, we show how our method generalizes to the Tecnick and CLIC dataset, followed by qualitative results. We analyze the stability of all functions and show the effect of interpolation between different methods with SSL. Lastly, we analyze a refining procedure at compression time that allows moving along the R-D curve when refining the latents with another $\lambda$ than a pre-trained model is trained on Gao et al. (2022); Xu et al. (2023). The code can be retrieved from: https://github.com/yperugachidiaz/flexible_neural_image_compression.
@@ 2 Preliminaries and related work
In lossy compression, the aim is to find a mapping of image $x$ where the distortion of the reconstructed image $\hat{x}$ is as little as possible compared to the original one while using as little storage as possible. Therefore, training a lossy neural image compression model presents a trade-off between minimizing the length of the bitstream for an image and minimizing the distortion of the reconstructed image Ballé et al. (2017); Lee et al. (2019); Minnen et al. (2018); Theis et al. (2017).
Neural image compression models from Ballé et al. (2017); Cheng et al. (2020); Minnen et al. (2018); Theis et al. (2017), also known as hyperpriors, accomplish this kind of mapping with latent variables. An image $x$ is encoded onto a latent representation $y=g_{a}(x)$, where $g_{a}(\cdot)$ is the encoder. Next, $y$ is quantized $Q(y)=\hat{y}$ into a discrete variable that is sent losslessly to the decoder. The reconstructed image is given by: $\hat{x}=g_{s}(\hat{y})$, where $g_{s}(\cdot)$ represents the decoder. The rate-distortion objective that needs to be minimized for this specific problem is given by:
@@ 2.1 Latent optimization
Neural image compression models have been trained over a huge set of images to find an optimal encoding. Yet, due to difficulties in optimization or due to constraints on the model capacity, model performance is sub-optimal. To overcome these issues, another type of optimizing compression performance is proposed in Campos et al. (2019); Yang et al. (2020) where they show how to find better compression results by utilizing pre-trained networks and keeping the encoder and decoder fixed but only adapting the latents. In these methods, a latent variable $y$ is iteratively adapted using differentiable operations at test time. The aim is to find a more optimal discrete latent representation $\hat{y}$. Therefore, the following minimization problem needs to be solved for an image $x$:
@@ 2.2 Stochastic Gumbel Annealing
$logits=(-\operatorname{atanh}(v_{L})/\tau,-\operatorname{atanh}(v_{R})/\tau).$ (3)
To obtain probabilities a softmax is used over the $logits$, which gives the probability $p(y)$ which is the chance of $v$ being floored: $p(y=\lfloor v\rfloor)$, or ceiled: $p(y=\lceil v\rceil)$. This is approximated by the Gumbel-softmax distribution. Then, samples are drawn: $\mathbf{y}\sim\text{Gumbel-Softmax}(logits,\tau)$ Jang et al. (2016) and are multiplied and summed with the vector $\mathbf{v}_{r}$ to obtain the quantized representation: $\hat{v}=\sum_{i}{(v_{r,i}*y_{i})}$. As SGA aids the discretization gap, this method may not have optimal performance and may not be as robust to changes in its temperature rate $\tau$.
Besides SGA, Yang et al. (2020) propose deterministic annealing Agustsson et al. (2017), which follows almost the same procedure as SGA, but instead of sampling stochastically from the Gumbel Softmax, this method uses a deterministic approach by computing probabilities with the Softmax from the $logits$. In practice, this method has been shown to suffer from unstable optimization behavior.
(a) Probability space for several SGA+ methods, along with $\operatorname{atanh}$. Solid lines denote the probability of flooring $\lfloor v\rfloor$ and dotted lines the probability of ceiling $\lceil v\rceil$.
(b) Three-class rounding for the extended version of linear. Solid lines denote two-class rounding, dashed lines denote three-class rounding and dotted lines indicate the smoothness.
@@ 2.3 Other methods
While methods such as SGA aim to optimize the latent variables for neural image compression at inference time, other approaches have been explored in recent research. Guo et al. (2021) proposed a soft-then-hard strategy alongside a learned scaling factor for the uniform noise to achieve better compression and a smoother latent. These methods are used to fine-tune network parameters but not the latents directly. Zhu et al. (2022) proposed using Swin-transformer-based coding instead of ConvNet-based coding. They showed that these transforms can achieve better compression with fewer parameters and shorter decoding times. van Rozendaal et al. (2021) proposed to also fine-tune the decoder alongside the latent for video compression. While accommodating the additional cost of saving the model update, they demonstrated a gain of $\sim 1dB$. Zhang et al. (2021) and Dupont et al. (2021) proposed using implicit neural representations for video and image compression, respectively. He et al. (2022) proposed an improved context model (SCCTX) and a change to the main transform (ELIC) that achieve strong compression results together. El-Nouby et al. (2023) revisited vector quantization for neural image compression and demonstrated it performs on par with hyperprior-based methods. Li et al. (2020) proposed a method to incorporate trellis-coded quantization in neural codecs. While these approaches change the training process, our work differs in that we only consider the inference process. Balcilar et al. (2023) proposes latent shift, a method that can further optimize latents using the correlation between the gradient of the reconstruction error and the gradient of the entropy.
@@ 3 Methods
As literature has shown, refining the latents of pre-trained compression models with SGA leads to improved compression performance Yang et al. (2020). In this section, we extend SGA by introducing SGA+ containing three other methods for the computation of the unnormalized log probabilities ($logits$) to overcome issues from its predecessor. We show how these methods behave in probability space. Furthermore, we show how the methods can be extended to three-class rounding.
@@ 3.1 Two-class rounding
Recall from SGA that a variable $v$ is quantized to indicate the rounding direction to two classes and is centered between (0,1). Computation of the unnormalized log probabilities is obtained with $\operatorname{atanh}$ from Equation (3). Recall, that in general the probabilities are given by a softmax over the logits with a function of choice. As an example, for SGA the logits are computed with $\operatorname{atanh}$. The corresponding probabilities for rounding down is then equal to: $\frac{e^{\operatorname{atanh}(v_{L})}}{e^{\operatorname{atanh}(v_{L})}+e^{\operatorname{atanh}(v_{R})}}$. Then looking at the probability space from this function, see Figure 1(a), the $\operatorname{atanh}$ function can lead to sub-optimal performance when used to determine rounding probabilities. The problem is that gradients tend to infinity when the function approaches the limits of 00 and $1$, see Appendix A for the proof that gradients at 0 tend to $\infty$. This is not ideal, as these limits are usually achieved when the discretization gap is minimal. In addition, the gradients may become larger towards the end of optimization. Further analyzing the probability space, we find that there are a lot of possibilities in choosing probabilities for rounding to two classes. However, there are some constraints: the probabilities need to be monotonic functions, and the probabilities for rounding down (flooring) and up (ceiling) need to sum up to one. Therefore, we introduce SGA+ and propose three methods that satisfy the above constraints and can be used to overcome the sub-optimality that the $\operatorname{atanh}$ function suffers from. We opted for these three as they each have their own interesting characteristics. However, there are many other functions that are also valid and would behave similarly to these three.
@@
where $a$ is the factor determining the shape of the function. SSL is exactly the linear for $a=1$. For $a=1.6$ and $a=0.65$ SSL roughly resembles the cosine and $\operatorname{atanh}$. For $a\rightarrow\infty$ the function tends to shape to (reversed) rounding.
Note that the main reason behind the linear version is the fact that it is the only function with constant gradients which is also the most robust choice, the cosine version is approximately mirrored across the diagonal of the $-\operatorname{atanh}(x)$ which shows that it is more stable compared to the $-\operatorname{atanh}(x)$, and the reason behind the SSL is that it is a function that can interpolate between all possible functions and can be tuned to find the best possible performance when necessary.
@@ 3.2 Three-class rounding
The probability that $v$ is rounded is denoted by: $p(y=\lfloor v\rceil)\propto f_{3c}(w|r,n)$, where $w=v-\lfloor v\rceil$ is centered around zero. Further, we define the probability that $v$ is rounded $+1$ and rounded $-1$ is respectively given by: $p(y=\lfloor v\rceil-1)\propto f_{3c}(w-1|r,n)$ and $p(y=\lfloor v\rceil+1)\propto f_{3c}(w+1|r,n)$. Recall, that $v_{L},v_{R}\in[0,1]$, whereas $w\in[-0.5,0.5]$. Defining $w$ like this is more helpful for the 3-class since it has a center class. The general notation for the three-class functions is given by:
@@
where $\operatorname{clip}(\cdot)$ clips the value at 00 and $1$, $r$ is the factor determining the height and steepness of the function and power $n$ controls the peakedness of the function. Note that $n$ can be fused with temperature $\tau$ together, to scale the function. This only accounts for the computation of the logits and not to modify the Gumbel temperature, therefore, $\tau$ still needs a separate definition.
@@
In Figure 1(b), three-class rounding for the extension of Equation (5) can be found. As can be seen, solid lines denote the special case of two-class rounding with $r=1$ and $n=1$, dashed lines denote three-class rounding with $r=0.9$ and $n=1$ and dotted lines denote the two-class rounding with $r=1$ and $n=3$, which shows a less peaked function. For an example of two- versus three-class rounding, consider the case where we have variable $v=-0.95$. For two-class rounding there is only the chance of rounding to $-1$ with $p(y=\lfloor v\rceil)$ (red solid line), a chance to round to 00 with $p(y=\lfloor v\rceil+1)$ (green solid line) and zero chance to round to $-2$ with $p(y=\lfloor v\rceil-1)$ (yellow solid line). For three-class rounding, with $r=0.9$ and $n=1$, when $v=-0.95$ we find a high chance to round to $-1$ with $p(y=\lfloor v\rceil)$ (red dashed line) and a small chance to round to 00 with $p(y=\lfloor v\rceil+1)$ (green dashed line) and a tiny chance to round to $-2$ with $p(y=\lfloor v\rceil-1)$ (yellow dashed line).
(a) True R-D loss
(b) Difference loss
(c) PSNR
(d) BPP
Figure 2: Performance plots of (a) True R-D Loss (b) Difference in loss (c) PSNR (d) BPP.